If n=CosA/CosB m=SinA/SinB So prove that- (m2-n2)Sin^2B=1-n 2

If n=CosA/CosB
    m=SinA/SinB
So prove that- 
(m2-n2)Sin^2B=1-n2
 

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
wehave to prove (m2-n2)Sin^2B=1-n
or equivalently m^2sin^B+n^2cos^2B=1
given that 
n=CosA/CosB and m=SinA/SinB
so, LHS = sin^2A+cos^2A=1=RHS 
hence proved